OpenStudy (anonymous):

I NEED HELP FINDIND THE COORDINATES?? GRAPH THE FOLLOWING LINE!!!! 7X-7Y=0 !!FIND TWO POINTS ON THE LINE TO BE GRAPHED AND ENTER THE COORDINATES A( , ) B( , )

OpenStudy (anonymous):

It can be any two points so you can make up and x or y value, plug it in, and get the corresponding x or y value. so well plug 2 numbers in for y, say 2 and 3 first one -6x+9(2)=0 x=?

OpenStudy (anonymous):

SO WILL IT BE 0,0

OpenStudy (anonymous):

No..solve the equation for x

OpenStudy (anonymous):

Well actually (0,0) could be one point on the line. so well use that point and whatever point we get from this equation

OpenStudy (anonymous):

I'm confused can you tell me what your coordinates will be if u was doing this problem

OpenStudy (anonymous):

-6X+9Y=0 you can get any point you want. plug in a number for y like this -6x+9(2)=0 -6x+18=0 -6x=-18 x=3 so the coordinate would be (3,2)

OpenStudy (anonymous):

so what if I was doing -6x-5y=0 so will I do -6x+5(6)=0 -6x=30 x=5,2

OpenStudy (anonymous):

i mean the coordinates will be 6,2

OpenStudy (anonymous):

Close...whatever you plug in for y will be the y coordinate, whatever x equals will be the x coordinate. -6x=-30 x=5 you got that by plugging in 6 for y so the coordinate is (5,6)

OpenStudy (anonymous):

so for y it will be -6(2)-5y=0 -12=0 coordinates will be 2,0

OpenStudy (anonymous):

Nope. the -5y disappeared... you're plugging in 2 for x so.. -6(2)-5y=0 -12-5y=0 -5y=12 y=-12/5 so your coordinate will be (2, -12/5)

OpenStudy (anonymous):

can you give me another number to plug in with the same equation for -6x-5y=0 for y

OpenStudy (anonymous):

-6x-5y=0 -5y=6x \[y=-\frac{ 6 }{5 }x\] when x=5,\[y=-\frac{ 6 }{5 }*5=-6\] so point is (5,-6) Again when x=0, y=0 Point is (0,0) now draw a line through these points.

OpenStudy (anonymous):

so how would i do it if i was doing 7x-7y=0

OpenStudy (anonymous):

i'm really confused so if you was doing this problem what will your coordinates be for A (: ) B:( )